# Hidden heavy flavor tetraquarks in the Born-Oppenheimer approximation

### v1

@@ III.2.3 The $c\bar{c}s\bar{s}$ configuration
$R_{0}=6$ 4.060, 4.150 4.096, 4.130 4.115 4.201
$R_{0}=8$ 4.041, 4.130 4.077, 4.111 4.096 4.182
$R_{0}=10$ 4.032, 4.122 4.069, 4.103 4.088 4.174
$R_{0}=12$ 4.030, 4.120 4.066, 4.101 4.086 4.172
III.3 Hidden-bottom tetraquark states with $b\bar{b}q\bar{q}$, $b\bar{b}q\bar{s}$, and $b\bar{b}s\bar{s}$ configurations
The hidden bottom quark tetraquark states can be realized by converting the charm quarks of the hidden charm tetraquark states into bottom quarks. These hidden bottom quark tetraquark states can be classified into three configurations: $b\bar{b}q\bar{q}$, $b\bar{b}q\bar{s}$, and $b\bar{b}s\bar{s}$. The mass spectrum at size $R_{0}=10GeV^{-1}$ is presented in FIG. 3. The analytical approach is similar to that used for the hidden charm tetraquark states. Therefore, we will not discuss each state in detail, but instead focus on the two states observed experimentally: $Z_{b}(10610)$ and $Z_{b}(10650)$. These states have been observed in the decay channels $\Upsilon(nS)\pi$ and $h_{b}(mP)$, where $n=1,2,3$ and $m=1,2$. Their quantum numbers are likely $J^{P}=1^{+}$, but there is no information regarding their $C$-parity. In our calculations, we prefer these two states to be the tetraquark states $b\bar{b}u\bar{d}$ or $b\bar{b}u\bar{d}$ with quantum number assignment $1^{+-}$, as shown in FIG. 3(a). In addition, we have calculated the masses of hidden-bottom tetraquark states at different sizes, as illustrated in Table 7-9, and the analysis method remains consistent.
Table 7: Mass spectra of tetraquark $b\bar{b}q\bar{q}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{++})$ $M(1^{+-})$ $M(1^{++})$ $M(2^{++})$
$R_{0}=6$ 10.603, 10.458 10.598, 10.455 10.605 10.621
$R_{0}=8$ 10.596, 10.451 10.591, 10.449 10.598 10.614
$R_{0}=10$ 10.593, 10.448 10.588, 10.446 10.595 10.611
$R_{0}=12$ 10.592, 10.447 10.587, 10.445 10.584 10.610
Table 8: Mass spectra of tetraquark $b\bar{b}q\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{+})$ $M(1^{+})$ $M(2^{+})$
$R_{0}=6$ 10.604, 10.694 10.604, 10.681, 10.720 10.684
$R_{0}=8$ 10.597, 10.687 10.598, 10.675, 10.713 10.678
$R_{0}=10$ 10.596, 10.685 10.596, 10.673, 10.712 10.676
$R_{0}=12$ 10.595, 10.685 10.595, 10.672, 10.711 10.676
Table 9: Mass spectra of tetraquark $b\bar{b}s\bar{s}$ states at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
$M(0^{++})$ $M(1^{+-})$ $M(1^{++})$ $M(2^{++})$
$R_{0}=6$ 10.696, 10.830 10.756, 10.834 10.789 10.877
$R_{0}=8$ 10.690, 10.824 10.750, 10.828 10.783 10.871
$R_{0}=10$ 10.688, 10.822 10.748, 10.826 10.781 10.869
$R_{0}=12$ 10.687, 10.821 10.747, 10.825 10.780 10.868
@@
In this paper, we investigate the hidden heavy flavor tetraquark states $Q\bar{Q}q\bar{q}$ composed of two heavy quarks and two light quarks. A remarkable feature of these tetraquark states is that they conform to a hydrogen-bond-like structure in quantum chromodynamics vertion, which can be treated using the BO approximation. Subsequently, we systematically calculated the mass spectra of $S$-wave tetraquark states $Q\bar{Q}q\bar{q}$ with various quantum numbers ($J^{P}=0^{+(+)}$, $1^{+(\pm)}$, and $2^{+(+)}$), and analyzed their decay channels.
@@
Our calculation results indicate that the exotic hadrons $X(4140)$, $Y(3940)$, $Z_{b}(10610)$, and $Z_{b}(10650)$ remain strong candidates for tetraquark states. Besides, we propose that $Z_{cs}(3985)^{-}$, $Z_{cs}(4000)^{+}$, and $Z_{cs}(4123)^{-}$ may be partner states of the tetraquark state $cu\bar{c}\bar{s}/cs\bar{c}\bar{u}$ with quantum number $J^{P}=1^{+}$. The $X(3960)$ may be a tetraquark state $cs\bar{c}\bar{s}$ with larger spatial sizes and quantum numbers $J^{PC}=1^{++}$. In addition, we predict several other possible hidden heavy flavor tetraquark states, and look forward to further experimental and theoretical research. Finally, it is worth emphasizing that although the difference in quark masses may have some impact on the mass spectra of tetraquark states, this hydrogen-bond-like structure in quantum chromodynamics should not be overlooked.

### v2

@@ III.2.3 The $c\bar{c}s\bar{s}$ configuration
$R_{0}=6$ 10799, 10764 10772, 10818 10819 10846
$R_{0}=8$ 10793, 10758 10766, 10812 10813 10840
$R_{0}=10$ 10791, 10756 10765, 10810 10812 10839
$R_{0}=12$ 10791, 10756 10764, 10810 10811 10838
The hidden-bottom tetraquark states can be realized by converting the charm quarks of the hidden-charm tetraquark states into bottom quarks. And these hidden-bottom tetraquark states can be classified into three configurations: $b\bar{b}n\bar{n}$, $b\bar{b}n\bar{s}$, and $b\bar{b}s\bar{s}$. The possible quantum number assignments for these $S$-wave tetraquark states are the same as for the hidden-charm tetraquark states, i.e., $J^{P(C)}=0^{+(+)}$, $1^{+(\pm)}$, and $2^{+(+)}$. Similarly, given $R_{0}=10$ GeV-1, we calculated the mass spectra of these states and plotted them in Figure 3. In addition, considering the influence of different $R_{0}$ on the mass spectra of hidden-bottom tetraquark states, we take $R_{0}=6,8,10$, and $12$ GeV-1 respectively, and the corresponding calculated mass spectra listed in Tables 6-8. The analytical approach for these states is similar to that used for the hidden-charm tetraquark states. Therefore, we will not discuss each state in detail.
In the subsection, we focus on the two exotic hadrons observed experimentally: $Z_{b}(10610)$ and $Z_{b}(10650)$. They are found in the decay channels $\pi^{\pm}\Upsilon(nS)$ and $\pi^{\pm}h_{b}(mP)$, where $n=1,2,3$ and $m=1,2$. This suggests that the component configuration of these two exotic hadrons may be a tetraquark state $b\bar{b}n\bar{n}$. And their quantum numbers are likely $J^{P}=1^{+}$, but there is no information regarding their $C$-parity. In our calculations, we prefer to interpret these two exotic states as the tetraquark states $b\bar{b}d\bar{u}$ or $b\bar{b}u\bar{d}$ with quantum number assignment $I(J^{PC})=1(1^{+-})$, as shown in Figure 3(a). However, the calculated masses of the tetraquark states are smaller than the corresponding experimental values for the two exotic hadrons. One obvious reason is that we chose a larger parameter $R_{0}$, see Table 6. A deeper reason is that the two quarks inside a hadron form a diquark pair $Qq$, which manifests as an attractive interaction. Compared with the hidden charm tetraquark system, the parameter $R_{0}$ corresponding to the hidden bottom tetraquark system should be smaller.
@@
In this paper, we investigate the hidden heavy flavor tetraquark states $Q\bar{Q}q\bar{q}$ composed of two heavy quarks and two light quarks. A remarkable feature of these tetraquark states is that they conform to a hydrogen-bond-like structure in quantum chromodynamics vertion, which can be treated using the BO approximation. Subsequently, we systematically calculated the mass spectra of $S$-wave tetraquark $Q\bar{Q}q\bar{q}$ system with various quantum numbers ($J^{P}=0^{+(+)}$, $1^{+(\pm)}$, and $2^{+(+)}$), and analyzed their decay channels.
@@
Our calculation results indicate that the exotic hadrons $X(4140)$, $Y(3940)$, $Z_{b}(10610)$, and $Z_{b}(10650)$ remain strong candidates for tetraquark states. Besides, we propose that $Z_{cs}(3985)^{-}$, $Z_{cs}(4000)^{+}$, and $Z_{cs}(4123)^{-}$ may be partner states of the tetraquark state $cu\bar{c}\bar{s}/cs\bar{c}\bar{u}$ with quantum number $J^{P}=1^{+}$. The $X(3960)$ may be a tetraquark state $cs\bar{c}\bar{s}$ with quantum number $J^{PC}=0^{++}$. In addition, we predict several other possible hidden heavy flavor tetraquark states, and look forward to further experimental and theoretical research. Finally, it is worth emphasizing that although the difference in quark masses may have some impact on the mass spectra of tetraquark states, this hydrogen-bond-like structure in quantum chromodynamics should not be overlooked.